how many grams of li are needed to produce 9.89 grams of h2


1
Expert's answer
2021-09-12T00:42:25-0400

Li + H2O --> LiOH + H2

Mole ratio = 1:1

Molar mass of H2 = 2.016

= 9.89/2.016 = 4.91 moles

Molar mass of Li = 6.941

= 6.941 × 4.91 = 34.08g